please help y=1/3+1 y=3x+3
Answer by Alan3354(69443)

please help y=1/3+1 y=3x+3
=========
y=1/3+1
y = 4/3
============
y=3x+3
4/3 = 3x + 3
3x = (4/3) - 3 = -5/3
x = -5/9